### v2

@@ Abstract
This paper proposes a correlated random coefficient linear panel data model, where regressors can be correlated with time-varying and individual-specific random coefficients through both a fixed effect and a time-varying random shock. I develop a new panel data-based identification method to identify the average partial effect and the local average response function. The identification strategy employs a sufficient statistic to control for the fixed effect and a conditional control variable for the random shock. Conditional on these two controls, the residual variation in the regressors is driven solely by the exogenous instrumental variables, and thus can be exploited to identify the parameters of interest. The constructive identification analysis leads to three-step series estimators, for which I establish rates of convergence and asymptotic normality. To illustrate the method, I estimate a heterogeneous Cobb-Douglas production function for manufacturing firms in China, finding substantial variations in output elasticities across firms.
Keywords: Correlated random coefficients, panel data, time-varying endogeneity, index exclusion, conditional control variable, average partial effect, local average response function.
@@ 1 Introduction
Correlated random coefficient (CRC) linear panel models have proven useful due to their ability to accommodate complex forms of unobserved heterogeneity that are empirically relevant (Wooldridge (2010), Hsiao (2022)). A crucial consideration in these models is the correlation between random coefficients and regressors. Classical methods typically address this issue by allowing a time-invariant, individual-specific fixed effect—either in the form of an additive intercept or an individual-specific coefficient—to be correlated with the regressors (e.g., Hausman and Taylor (1981), Hsiao and Pesaran (2008)). While convenient, this approach may not fully capture agents’ optimization behavior. For instance, it is possible that a firm with a high fixed effect (e.g., strong management skills) endogenously selects lower input levels during certain periods. Such behavior could arise if the firm has information about its time-varying output efficiency shock, in addition to its time-invariant management skill, when optimally determining input values.
In this paper, I aim to address this gap by proposing a new time-varying endogenous random coefficient (TERC) linear panel model, where regressors can be correlated with time-varying and individual-specific random coefficients not only through a fixed effect but, more importantly, through a time-varying random shock. Specifically, the baseline TERC model [^1] In the baseline model, I focus exclusively on endogenous regressors to address the issue of time-varying endogeneity through the random coefficients. The analysis remains largely unchanged when exogenous regressors are included in (1.1). I specify the necessary changes at the end of the proof of Theorem 1 in the Appendix. consists of two equations:
$Y_{it}$ $=X_{it}^{{}^{\prime}}\beta\left(A_{i},\varepsilon_{it}\right),$ (1.1)
@@
In (1.1), I model the vector of random coefficients $\beta_{it}=\beta\left(A_{i},\varepsilon_{it}\right)$ as a vector of unknown and possibly nonlinear functions $\beta\left(\cdot\right)$ of a fixed effect $A_{i}$ and a time-varying random shock $\varepsilon_{it}$. $Y_{it}$ is then determined by the inner product between $X_{it}$ and $\beta_{it}$. In (1.2), I model the vector of regressors $X_{it}$ as a vector-valued, unknown and possibly nonlinear function $g\left(\cdot\right)$ of the vector of exogenous instrumental variables (IV) $Z_{it}$, fixed effect $A_{i}$, and per-period information $\eta_{it}$ about $\varepsilon_{it}$. The motivation for (1.2) is that the agent $i$ in period $t$ optimally chooses values of her regressors $X_{it}$ by solving an optimization problem (e.g., firm profit maximization), with $Z_{it},\ A_{i},\text{ and }\eta_{it}$ in her information set, leading to (1.2). All of $A_{i}$, $\varepsilon_{it}$, and $\eta_{it}$ can be scalar- or vector-valued [^2] In the baseline model, I assume in Assumption 1 that each coordinate of $g\left(\cdot\right)$ depends monotonically on a single coordinate of $\eta_{it}$ to invoke the control function argument. As an extension, I discuss how to relax this assumption to allow the entire vector $\eta_{it}$ to enter certain coordinates of $g\left(\cdot\right)$ at the end of Section 3., and any pair of the three variables can be correlated. The functional forms are potentially different across different coordinates of $\beta\left(\cdot\right)$ (and also $g\left(\cdot\right)$). As an analyst, I observe $\left\{X_{it},Y_{it},Z_{it}\right\}$ for $i=1,...,n$ and $t=1,...,T$ and aim to identify the average partial effect $\mathbb{E}\beta_{it}$ (APE, see Graham and Powell (2012)) and the local average response function $\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]$ (LAR, see Altonji and Matzkin (2005)).
@@
A key feature of the TERC model is that $X_{it}$ can be correlated with $\beta_{it}$ through both $A_{i}$ and $\eta_{it}$, potentially in a complicated manner. I refer to this feature as “time-varying endogeneity through the random coefficients.” Allowing such correlation is important to applications when agent $i$ in period $t$ has information about $\beta_{it}$ through both $A_{i}$ and $\eta_{it}$ when optimally deciding $X_{it}$. I present three empirical examples in Section 2 that demonstrate the validity of such correlation in real-life scenarios. The associated technical challenge is that I need to control for the time-varying endogeneity through the random coefficients when both unknown (to the analyst) $A_{i}$ and $\eta_{it}$ appear in $g\left(\cdot\right)$ in a nonlinear and nonseparable way in (1.2). Furthermore, I do not specify the distributions of $A_{i}$, $\varepsilon_{it}$, or $\eta_{it}$, nor the functional form of $\beta\left(\cdot\right)$ or $g\left(\cdot\right)$. In this sense, the TERC model is a fixed effect panel data model (Laage (2024)). It has many important applications in economics; e.g., heterogeneous Cobb-Douglas production function estimation, return to schooling estimation, labor supply estimation, Engel curve analysis, and demand analysis, among others (Blundell, MaCurdy, and Meghir (2007b, a), Wooldridge (2009), Chernozhukov, Hausman, and Newey (2019), Li and Sasaki (2024), Keiller, de Paula, and Van Reenen (2024)).
[…]
The constructive identification analysis leads to easy-to-implement series estimators for the APE and LAR. I derive the convergence rates and establish the asymptotic normality of the proposed estimators. The inference results presented in this paper build on the literature on the asymptotic analysis of multi-step series estimators (Andrews (1991), Newey (1997), Imbens and Newey (2009)). The key departures from the existing literature include the fact that the object of interest is a partial mean process (Newey (1994)) of the derivative of the second-step estimator, and that the final step of the three-step estimator is an unknown but estimable functional of a conditional expectation function. Consequently, the estimation errors from each step must be properly accounted for to ensure accurate large sample properties.
As an empirical illustration, I estimate a Cobb-Douglas production function with heterogeneous output elasticities for Chinese manufacturing firms. My estimated APEs are broadly consistent with those derived from applying the classical methods with constant coefficients (Olley and Pakes (1996), Levinsohn and Petrin (2003), Ackerberg, Caves, and Frazer (2015)) to the same data. I also estimate the LARs evaluated at the realized input levels for each firm and find significant variation across firms. [^4] Note that, with constant coefficient models, the LAR is equivalent to the APE. To support the empirical results, I conduct simulations motivated by production function estimation applications to evaluate the finite-sample performance of the estimators proposed in this paper. The results demonstrate that the proposed method performs well.
Next, I discuss the contributions and limitations of this paper. The first contribution concerns the features of the TERC model in (1.1)–(1.2). Allowing $X_{it}$ to be correlated with $\beta_{it}$ not only through $A_{i}$ but, more importantly, through $\eta_{it}$ significantly enhances the model’s flexibility and carries important empirical implications, as demonstrated by the examples discussed so far. Another novel aspect of the TERC model is the inclusion of $A_{i}$ in both (1.1) and (1.2). In particular, I argue that including $A_{i}$ in the first-step function $g\left(\cdot\right)$ better reflects the reality that $A_{i}$ is typically in the information set of agent $i$ when she optimally chooses the value of $X_{it}$, for instance, through repeated learning. Moreover, $A_{i}$ enters both (1.1) and (1.2) in a nonlinear and nonseparable way, which naturally arises from agent optimization behavior. Consequently, enabling these features enhances the model’s flexibility and broadens its applicability.
The second contribution is the proposal of a new panel data-based method to identify the APE and LAR of the TERC model. I construct a sufficient statistic $W_{i}$ to control for the time-invariant $A_{i}$ and a per-period conditional control variable $V_{it}$ to control for the time-varying shock $\eta_{it}$. I prove that, conditional on $V_{it}\text{ and }W_{i}$, the residual variation in $X_{it}$ is entirely attributable to the exogenous $Z_{it}$, thereby enabling the identification of the APE and LAR. The identification method proposed in this paper also contributes to the literature on control functions with multiple unobservables in nonseparable regression models (e.g., Blundell, Kristensen, and Matzkin (2013), Kasy (2014)).
The final contribution involves implementing the proposed method and characterizing its statistical properties. I develop three-step series estimators for the APE and LAR, establishing their convergence rates and asymptotic normality. The effectiveness of the method is demonstrated using a production dataset of Chinese manufacturing firms, revealing significant heterogeneity in capital and labor elasticities across firms. To further support the empirical findings, I conduct a simulation study inspired by applications in production function estimation, which produces satisfactory results.
I discuss two limitations of this paper. First, although both the unknown $A_{i}$ and $\eta_{it}$ determine the value of $X_{it}$, the baseline model imposes a restriction such that each coordinate of $g\left(\cdot\right)$ depends on a single coordinate of $\eta_{it}$ (as well as $Z_{it}\text{ and }A_{i}$). As noted by Imbens and Newey (2009), this restriction excludes nonseparable supply and demand models with one time-varying disturbance per equation. To relax this restriction, I propose two approaches under additional empirically motivated assumptions at the end of Section 3. Second, my method requires IVs. I provide a few examples of IVs for the applications discussed in this paper. However, depending on the specific application, identifying suitable IVs can still be challenging.
@@ 1.1 Related Literature
This paper contributes to the literature on CRC linear panel models. I highlight a few studies that are particularly relevant to this work. Chamberlain (1992) studies the regular identification of the first-order moments of the random coefficients when $T>d_{X}$, where $T$ is the number of periods and $d_{X}$ is the dimension of regressors. Wooldridge (2005) derives conditions for consistency of the fixed-effects estimators of the APE for a class of linear and nonlinear panel data models with time-invariant random coefficients. The condition requires the random coefficients to be mean independent of the residuals from regressing the covariates on the aggregate time variables. Graham and Powell (2012) prove irregular identification when $T=d_{X}$ and allow for more persistent processes in the regressors. They leverage different identifying content of the subpopulations of “movers” and “stayers” to identify the first-order moments of the time-varying random coefficients. Arellano and Bonhomme (2012) identify the first-order moments of the time-invariant random coefficients by using the within-group and between-group variation in the regressors. They rely on a mean independence assumption between the residuals and both the regressors and the random coefficients. They further identify the second-order moments and distribution functions of the random coefficients by exploiting assumptions on the time-dependence structure of the residuals. Graham, Hahn, Poirier, and Powell (2018) show how to identify quantiles of the random coefficients in a general linear quantile regression model. Masten (2018) considers a classical linear simultaneous equations model with random coefficients on the endogenous variables. By extending the Chamberlain-Mundlak approach for balanced panels, Wooldridge (2019) develops estimation strategies for CRC linear panel models that account for sample selection in unbalanced panels. Laage (2024) studies a time-invariant CRC linear panel data model with additive scalar-valued fixed effects. She proposes a two-step identification approach, where in the first step she identifies nonparametrically the conditional expectation of the disturbances given the regressors and the control variables, and then in the second step she uses between-group variation to identify the APE.
The key difference between my paper and the CRC literature discussed above is that my model, (1.1)–(1.2), features time-varying random coefficients that can be correlated with the regressors through both fixed effects and time-varying shocks. None of the papers, except Graham and Powell (2012), explicitly considers such correlation. Graham and Powell (2012) impose a time-stationarity assumption on the time-varying random shocks given the regressors for all periods and the fixed effect, and exploit the different identifying contents of the subpopulations for identification. The novel method proposed by Graham and Powell (2012) covers a wide range of empirical applications. However, their main time-stationarity assumption effectively rules out the time-varying endogeneity through the random coefficients, wherein $X_{it}$ and $\beta_{it}$ are correlated through both $A_{i}$ and $\eta_{it}$ [^5] To be precise, the time-stationarity assumption imposed by Graham and Powell (2012) is $\left.\eta_{it}\sim_{d}\eta_{is}\right|{\bf X}_{i},A_{i},\text{ for }t\neq s.$ To see why it rules out time-varying endogeneity through the random coefficients, consider a simplified example in which the number of periods $T=2$ and the data generating process (DGP) of a scalar-valued $X_{it}$ is $X_{it}=Z_{it}^{{}^{\prime}}\lambda+A_{i}+\eta_{it}$. Suppose one observes $X_{i2}>X_{i1}$ and $Z_{i2}=Z_{i1}$ in the data, which by the DGP of $X_{it}$ implies $\eta_{i2}>\eta_{i1}$ deterministically given ${\bf X}_{i}$ and $A_{i}$. Consequently, the time-stationarity condition is violated when the time-varying endogeneity through the random coefficients is present., which is the focus of this paper. Instead, I tackle this issue by proposing two controls for $A_{i}$ and $\eta_{it}$ and, subsequently, using the residual variation in $X_{it}$ to identify the APE and LAR. It should be clarified that my work is not necessarily more general than that of Graham and Powell (2012), as the set of assumptions used by each is not nested within the other’s. Instead, I consider this paper an initial step toward addressing the important issue of time-varying endogeneity through the random coefficients, which existing methods have not yet fully resolved.
@@
The second line of research concerns the techniques used in this paper. My method requires the existence of a sufficient statistic $W_{i}$ for the time-invariant $A_{i}$ and a feasible control variable $V_{it}$ for the time-varying $\eta_{it}$. To justify the sufficiency of $W_{i}$ for $A_{i}$, I present several techniques from the literature (e.g., Mundlak (1978), Altonji and Matzkin (2005), Bester and Hansen (2009), Wooldridge (2019), Arkhangelsky and Imbens (2022)). See Liu, Poirier, and Shiu (2024) for a detailed discussion on the index sufficiency condition for the fixed effects in nonlinear semiparametric panel data models. It may be worth pointing out that the exchangeability condition used by Altonji and Matzkin (2005) concerns the conditional probability density function (PDF) of $A_{i}\text{ and }\eta_{it}$ given ${\bf X}_{i}$. I show that it is not directly applicable to the TERC model in which $X_{it}$ depends on $\eta_{it}$ for each $t$. Instead, I impose a lower-level exchangeability condition involving only the unobservable $A_{i}$ and $\eta_{it}$ and use it to establish a sufficiency result for $A_{i}$ in Proposition 1.
To find the feasible control variable for $\eta_{it}$ in (1.2), I generalize the technique of Imbens and Newey (2009) in two nontrivial ways. First, because the unknown $A_{i}$ enters equation (1.2) together with $\eta_{it}$, the conditional CDF $F_{\left.X_{it}\right|Z_{it}}\left(\left.X_{it}\right|Z_{it}\right)$ proposed by Imbens and Newey (2009) does not uniquely determine $\eta_{it}$ even if it is a scalar. Instead, I use the index exclusion condition on $A_{i}$ to construct a feasible conditional control variable $V_{it}=F_{\left.X_{it}\right|Z_{it},W_{i}}\left(\left.X_{it}\right|Z_{it},W_{i}\right)$ and show that it controls for $\eta_{it}$ given $A_{i}$ and $W_{i}$. Second, I allow $\eta_{it}$ to be vector-valued, enabling the model to capture the rich heterogeneity commonly observed in empirical applications. In the baseline model, I assume that each coordinate of $X_{it}$ depends upon a single coordinate of $\eta_{it}$. As an extension, I discuss how to relax this restriction when additional empirically motivated assumptions are imposed. Moreover, constructing $V_{it}$ requires IVs that satisfy specific exogeneity conditions. Murtazashvili and Wooldridge (2008) provide a set of conditions sufficient for the consistency of a general class of fixed effects IV estimators in a time-invariant individual-specific CRC panel data model. The model and method of this paper are different from theirs.
@@
The asymptotic analysis in this paper builds on a series of foundational works. Andrews (1991) investigates the asymptotic properties of series estimators for nonparametric and semiparametric regression models. His results are applicable to a wide variety of estimands, including derivatives and integrals of the regression function. I use his results to cover vector-valued functionals of regression functions. Newey (1997) also studies series estimators and provides conditions for proving convergence rates and asymptotic normality for the estimators of conditional expectations. Newey, Powell, and Vella (1999) propose a two-step nonparametric estimator for a triangular simultaneous equation model with a separable first-step equation. They derive asymptotic normality for their two-step estimator with the first-step estimator plugged in. Imbens and Newey (2009) also analyze a triangular simultaneous equation model, but with a nonseparable first-step equation. They obtain mean-squared convergence rates for the first-step estimator and prove asymptotic normality (in their 2002 working paper version) for known functionals of the conditional expectation of the outcome variable given the regressors and control variables. I extend their analysis to establish asymptotic normality for unknown but estimable functionals of conditional expectation functions.
@@
The rest of the paper is organized as follows. Section 2 formally introduces the TERC model and provides several empirical examples that fit within its framework. Section 3 outlines the identification idea and presents the assumptions, key identification theorem, and three extensions. Section 4 presents the series estimators for the APE and LAR and establishes their asymptotic properties. Section 5 provides an empirical illustration. Finally, Section 6 concludes. Proofs of all theorems, Proposition 1, and a simulation study are included in Appendix A, B, and C, respectively.

### v3

@@ Abstract
This paper proposes a correlated random coefficient linear panel data model, where regressors can be correlated with time-varying and individual-specific random coefficients through both a fixed effect and a time-varying random shock. I develop a new panel data-based method to identify the average partial effect and the local average response function. The identification strategy employs a sufficient statistic to control for the fixed effect and a control variable for the random shock. Conditional on these two controls, the residual variation in the regressors is driven solely by the exogenous instrumental variables, and thus can be exploited to identify the parameters of interest. The constructive identification analysis leads to three-step series estimators, for which I establish rates of convergence and asymptotic normality. To illustrate the method, I estimate a heterogeneous Cobb-Douglas production function for manufacturing firms in China, finding substantial variations in output elasticities across firms that can be related to various firm characteristics.
Keywords: Correlated random coefficients, panel data, time-varying endogeneity, semiparametric estimation.
@@ 1 Introduction
Correlated random coefficient (CRC) linear panel models have proven useful due to their ability to accommodate complex forms of unobserved heterogeneity that are empirically relevant (Wooldridge (2010), Hsiao (2022)). A crucial consideration in these models is the correlation between random coefficients and regressors. Classical methods typically address this issue by allowing a time-invariant, individual-specific fixed effect—either in the form of an additive intercept or an individual-specific coefficient—to be correlated with the regressors (e.g., Hausman and Taylor (1981), Hsiao and Pesaran (2008)). While convenient, this approach may not fully capture agents’ optimization behavior. For example, even high-capability firms may cut inputs in some periods when input prices are flat because input choices also respond to transitory, time-varying shocks to output efficiency or productivity. Throughout the sample, I assume capability is time-invariant. Thus, both persistent capability and transitory shocks can affect realized output efficiency and may be reflected in input choices, thereby generating endogeneity that is not fully absorbed by a fixed effect.
In this paper, I aim to address this gap by proposing a new time-varying endogenous random coefficient (TERC) linear panel model, where regressors can be correlated with time-varying and individual-specific random coefficients not only through a fixed effect but, more importantly, through a time-varying random shock. Specifically, the baseline TERC model consists of two equations:
$Y_{it}$ $=X_{it}^{\prime}\beta\left(A_{i},\varepsilon_{it}\right),$ (1.1)
@@
The vector of random coefficients $\beta_{it}=\beta\left(A_{i},\varepsilon_{it}\right)$ is modeled as a vector of unknown functions $\beta$ of a fixed effect $A_{i}$ and a time-varying random shock $\varepsilon_{it}$. $Y_{it}$ is then determined by the inner product between $X_{it}$ and $\beta_{it}$. In (1.2), I model the vector of regressors $X_{it}$ as a vector-valued, time-varying function $g$ of the vector of exogenous instrumental variables (IV) $Z_{it}$, fixed effect $A_{i}$, and per-period information $\eta_{it}$ about $\beta_{it}$.
@@
The motivation for (1.2) is that the agent $i$ in period $t$ optimally chooses values of her regressors $X_{it}$ by solving an optimization problem (e.g., firm profit maximization), with $Z_{it},\ A_{i},\text{ and }\eta_{it}$ in her information set, leading to (1.2). All of $A_{i}$, $\varepsilon_{it}$, and $\eta_{it}$ can be scalar- or vector-valued, and any pair of the three variables can be correlated. As an analyst, I observe $\left\{X_{it},Y_{it},Z_{it}\right\}$ for $i=1,\ldots,n$ and $t=1,\ldots,T$ and aim to identify the pooled average partial effect $T^{-1}\sum_{t=1}^{T}\mathbb{E}\beta_{it}$ (APE, see Graham and Powell (2012)) and the local average response function $\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]$ (LAR, see Altonji and Matzkin (2005)). [^1] Note that the joint distribution of $(\varepsilon_{it},\eta_{it})$ may vary over $t$. Consequently, any functionals of their distribution (e.g., $\mathbb{E}\beta_{it}$ and $\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]$) are in principle time-varying. I suppress the subscript $t$ whenever it is clear from context.
@@
A key feature of the TERC model is that $X_{it}$ can be correlated with $\beta_{it}$ through both $A_{i}$ and $\eta_{it}$, potentially in a complicated manner. I refer to this feature as “time-varying endogeneity through the random coefficients.” Allowing such correlation is important to applications when agent $i$ in period $t$ has information about $\beta_{it}$ through both $A_{i}$ and $\eta_{it}$ when optimally deciding $X_{it}$. In Section 2, I present three empirical examples—heterogeneous production function estimation, a labor supply model, and demand estimation—that demonstrate the validity of this correlation (Deaton (1986), Blundell, MaCurdy, and Meghir (2007), Li and Sasaki (2024), Keiller, de Paula, and Van Reenen (2024)). The associated technical challenge is to control for the time-varying endogeneity through the random coefficients when both $A_{i}$ and $\eta_{it}$ enter $g$ in a nonlinear and nonseparable manner in (1.2). Moreover, I do not impose parametric assumptions on the distributions of $A_{i}$, $\varepsilon_{it}$, or $\eta_{it}$, nor do I specify the functional form of $\beta$ or $g$. In this sense, the TERC model is a fixed-effects panel data model, and addressing these challenges requires a new method.
I propose a new panel data-based method for identifying the APE and LAR in the TERC model. The idea is to control for $A_{i}$ and $\eta_{it}$ via a sufficient statistic and a generalized residual, respectively, such that, conditional on these two controls, the residual variation in $X_{it}$ is driven solely by the exogenous IV $Z_{it}$. In other words, the residual variation in $X_{it}$ is causal and thus can be exploited to identify the parameters of interest. Specifically, I first impose an index exclusion assumption that supplies a sufficient statistic $W_{i}$ for $A_{i}$. I present justifications for this assumption from the literature. Next, I construct a feasible conditional cumulative distribution function (CDF) $V_{it}=F_{\left.X_{it}\right|Z_{it},W_{i}}\left(\left.X_{it}\right|Z_{it},W_{i}\right)$ to control for $\eta_{it}$. Given $(A_{i},W_{i})$, I establish a one-to-one mapping between $V_{it}$ and $\eta_{it}$. Finally, conditioning on these two controls—which effectively fixes both $A_{i}$ and $\eta_{it}$—(1.2) implies that the residual variation in $X_{it}$ is driven solely by $Z_{it}$. This residual, instrument-induced variation can therefore be exploited to identify the APE and LAR via perturbation arguments and iterated expectations. At the end of Section 3, I present three extensions: (i) allowing $g$ to depend on multiple components of $\eta_{it}$, (ii) identifying higher-order moments of $\beta_{it}$, and (iii) adding ex-post shocks to $\beta_{it}$ and $Y_{it}$, and including exogenous covariates into $X_{it}$.
I estimate the parameters of interest using a three-step series procedure. In the first step, I estimate the conditional CDF $V_{it}$, which serves as a control for $\eta_{it}$. In the second step, I exploit the linear structure of (1.1) to estimate $\mathbb{E}[\left.Y_{it}\right|X_{it},V_{it},W_{i}]$. In the third step, I use these estimates to recover the APE and LAR. When working with large datasets, the main computational challenge comes from the first-step estimation of $V_{it}$ and the second-step estimation of $\mathbb{E}[\left.Y_{it}\right|X_{it},V_{it},W_{i}]$. I address these issues in Remark 3.
Inference is less standard because these estimators are multi-step: they rely on intermediate nonparametric estimates, include derivatives of an earlier-step estimate, and then apply an additional unknown but estimable mapping to a conditional expectation. As a result, sampling error from the first and second steps can affect the limiting distribution of the final estimators, so valid standard errors must account for how uncertainty propagates across steps. I therefore derive convergence rates and asymptotic normality by explicitly tracking the contribution of each step, building on existing results for multi-step series estimators (Andrews (1991), Newey (1997), Imbens and Newey (2009)). A discussion of how the errors from each step aggregate into the final standard error is provided at the end of Subsection 4.2.
To illustrate the method, I estimate a Cobb–Douglas production function with heterogeneous output elasticities using data on Chinese manufacturing firms. The instruments are inspired by Berry, Levinsohn, and Pakes (1995) and constructed as competitors’ leave-one-out weighted-average input prices within the same city and industry. The resulting APE estimates are broadly consistent with those obtained from classical constant-coefficient approaches applied to the same data (Olley and Pakes (1996), Levinsohn and Petrin (2003), Ackerberg, Caves, and Frazer (2015)). I also estimate the LAR functions evaluated at each firm’s realized input choices, and I find substantial cross-firm variation. I then relate heterogeneity in output elasticities to firm characteristics and obtain economically intuitive patterns. Finally, I conduct extensive robustness checks in Appendix B and find that the results are stable.
To complement the empirical analysis, Appendix C presents Monte Carlo simulations designed to mirror the empirical environment; the results show that the proposed estimator performs well and reinforce the empirical findings. In particular, in the baseline design ($n=1,000$, $T=2$, and 1,000 replications, using second-degree spline bases), the APE estimator exhibits small normalized bias (about 2.8–3.2% relative to the length of the true parameter) and low normalized rMSE (about 2.9–3.7%) across coefficients. Estimating the function-valued LAR is naturally less precise, but still yields normalized rMSEs on the order of 3.4–5.6% for all the coefficients. Performance improves with larger $n$ and additional time periods $T$. Robustness checks also indicate that second-degree splines outperform first- or third-degree bases, and that adding ex-post shocks increases rMSEs only slightly.
Before turning to the related literature, I note two limitations of the paper. First, the baseline model assumes each coordinate of $g$ is monotone in a single coordinate of $\eta_{it}$, which (as in Imbens and Newey (2009)) rules out certain nonseparable supply-and-demand systems. Subsection 3.3 discusses remedies under additional, empirically motivated assumptions. Second, the approach requires IVs and a tractable way to summarize the fixed effect $A_{i}$ via a sufficient statistic $W_{i}$; while I provide examples for the applications considered, finding credible instruments (and justifying the required structure) may be challenging in some settings.
@@ 1.1 Related Literature
This paper contributes to the literature on CRC linear panel models. Related work includes Chamberlain (1992) on regular identification of first moments when $T>d_{X}$; Graham and Powell (2012) on irregular identification when $T=d_{X}$ using movers/stayers; and Arellano and Bonhomme (2012) on identifying first (and, under additional assumptions, higher) moments and distributions via within–between variation and restrictions on residual dependence. Wooldridge (2005, 2009) develops estimation strategies for APEs—respectively under mean-independence conditions and with sample selection in unbalanced panels—while Murtazashvili and Wooldridge (2008) study consistency of fixed-effects IV estimators in time-invariant CRC models. Additional contributions include identification of quantiles (Graham, Hahn, Poirier, and Powell (2018)), random-coefficient simultaneous equations (Masten (2018)), and a two-step identification approach with additive fixed effects and time-invariant random coefficients (Laage (2024)).
My contribution to the CRC literature is to study a model, (1.1) and (1.2), with time-varying random coefficients that can be correlated with regressors through both fixed effects and time-varying shocks. The closest paper to my work is Graham and Powell (2012), who identify moments by exploiting movers and stayers under a time-stationarity restriction. This time-stationarity restriction conditions on the regressors for all periods and thus effectively rules out the time-varying endogeneity central here. I instead introduce controls for $A_{i}$ and importantly $\eta_{it}$ to identify the APE and LAR. The two approaches rely on non-nested assumptions; I view this paper as a step toward addressing time-varying endogeneity through random coefficients.
@@
The second line of research concerns the techniques used in this paper. For nonseparable panel models with generalized fixed effects, restrictions on the fixed effect are generally needed to identify global objects such as the APE and LAR (Hoderlein and White (2012, p.2); see also Gao, Li, and Xu (2023) and Gao and Li (2026) on dealing with the issue of fixed effects in network formation and panel multinomial choice models, respectively). My method requires a sufficient statistic $W_{i}$ for $A_{i}$ and a control variable $V_{it}$ for $\eta_{it}$. To justify the sufficiency of $W_{i}$ for $A_{i}$, I adapt several techniques from the literature (e.g., Mundlak (1978), Altonji and Matzkin (2005), Bester and Hansen (2009), Wooldridge (2019), Arkhangelsky and Imbens (2022)). See Liu, Poirier, and Shiu (2025) for a discussion on the index sufficiency condition in nonlinear semiparametric panel data models. To justify $V_{it}$ as a valid control for $\eta_{it}$, I generalize the technique of Imbens and Newey (2009) in two nontrivial ways and discuss them in relation to (3.15). More recently, Nagasawa (2024) studies treatment effect estimation with noisy conditioning variables using control variables from the joint distribution of noisy proxy measures.
@@
The asymptotic analysis builds on foundational results for series and control-function estimators. Andrews (1991) and Newey (1997) provide conditions for convergence rates and asymptotic normality of series estimators, which I use to handle vector-valued functionals of regression functions. Newey, Powell, and Vella (1999) establish asymptotic normality for two-step nonparametric estimators in triangular models with a separable first stage, while Imbens and Newey (2009) extend the triangular framework to nonseparable first stages and derive rates and asymptotic normality for known functionals of conditional expectations. I build on Imbens and Newey (2009) to obtain asymptotic normality for unknown but estimable functionals of conditional expectation functions.
@@
Organization. The rest of the paper is organized as follows. Section 2 formally introduces the TERC model and provides several empirical examples that fit within its framework. Section 3 outlines the identification idea and presents the assumptions, key identification theorem, and three extensions. Section 4 presents the series estimators for the APE and LAR and establishes their asymptotic properties. Section 5 provides an empirical illustration. Finally, Section 6 concludes. Proofs, additional empirical results, and a simulation study are included in Appendix A, B, and C, respectively.
Notation. Let $i\in\left\{1,\ldots,n\right\}$ index agents and $t\in\left\{1,\ldots,T\right\}$ index periods, with finite $T\geq 2$. I use boldface upper case letters for random matrices, regular upper case letters for random variables and random vectors, and lower case letters for their values. Let ${\bf X}_{i}=\left(X_{i1},...,X_{iT}\right)^{\prime}$, $Y_{i}=\left(Y_{i1},...,Y_{iT}\right)^{\prime}$, ${\bf Z}_{i}=\left(Z_{i1},...,Z_{iT}\right)^{\prime}$, $\varepsilon_{i}=\left(\varepsilon_{i1},...,\varepsilon_{iT}\right)^{\prime}$, and $\eta_{i}=\left(\eta_{i1},...,\eta_{iT}\right)^{\prime}$. I assume $\left({\bf X}_{i},{\bf Z}_{i},A_{i},\varepsilon_{i},\eta_{i}\right)$ are i.i.d. across $i$. Let $d_{X}$ be the dimension of $X_{it}$ and $X_{it,l}$ denote the $l^{\text{th}}$ coordinate of vector $X_{it}$. Similar notation is used for all other variables. I use $\coloneqq$ to define a new random variable, $\sim_{d}$ to indicate that two random variables are identically distributed, $F$ for CDF, $f$ for probability density function (PDF), $\mathbb{E}$ for expectation, $\mathbb{\mathbb{V}}$ for variance, $I_{d}$ for $d\times d$ identity matrix, $\overset{p}{\longrightarrow}$ for convergence in probability, and $\overset{d}{\longrightarrow}$ for convergence in distribution. All convergence results are stated under $n\to\infty$.